OpenStudy (anonymous):

solve log4 (3x-1) = log4(2x+3) (4 is the base)

OpenStudy (anonymous):

3x-1 = 2x+3 3x-2x =3+1 x=4

OpenStudy (kirbykirby):

If you have the same log base on both sides, you can just cancel them out :), which is what @matricked did

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

welcome